# A Study of Library Usage in Agent-Authored Pull Requests

### v1

@@ 4.2. Findings for RQ2 New Dependencies
For this RQ, we analyse dependency manifest files to investigate how often agents add new dependencies to projects. Full analysis is shown in Table 2.
Across all languages, dependency addition is extremely rare: only 1.3% of PRs introduce even a single new library. C# and Python stand out as the only languages where dependency addition is non-negligible (6.1% and 2.8% of PRs). This conservative behaviour contrasts with typical human workflows in large ecosystems—such as TypeScript and Python—where adding a dependency to reuse existing functionality is common practice (abdalkareemImpactUsingTrivial2020). Recent work shows that agents often struggle with installing packages and resolving environment conflicts (aroraSetupBenchAssessingSoftware2025), suggesting that their reluctance to add dependencies may reflect practical limitations rather than deliberate conservatism.
When agents do introduce new dependencies, they generally do so with good versioning practices. Across all languages, 75.0% of new dependencies include an explicit version constraint, with TypeScript and Go specifying versions in every instance. This contrasts with previous, non-agentic, LLM research that observes version constraints mentioned in only $\approx$9% of library-related discussions (rajRoleLibraryVersions2024a). Whilst not directly comparable, this does indicate that agentic setups may support more robust dependency management over direct LLM usage; likely because agents benefit from full project context and direct interaction with dependency manifests.
Finally, we observe substantial diversity in the dependencies added: 1,120 unique libraries across 2,129 total additions. This high ratio of unique packages suggests that agents do not default to a small, preferred set of libraries when adding new dependencies. Much of this diversity comes from a small number of diverse PRs that create entirely new projects and therefore introduce many dependencies at once; a pattern consistent with known strengths of modern LLMs in scaffolding initial project structures (rasnayakaEmpiricalStudyUsage2024).
RQ2 Summary
Agents rarely add new dependencies (1.3% of PRs), most commonly for C# (6.1%) and Python (2.8%). When they do add dependencies, they have strong version hygiene, specifying a constraint in 75.0% of cases.
@@ 4.3. Findings for RQ3 Choosing Libraries
For this RQ, we analyse trends in the libraries that agents most commonly import and add as new dependencies. Table 3 lists the most common libraries.
First, testing and development-tooling libraries appear across all languages. Although test frameworks are heavily-used, they make up only a minority of the distinct most-imported libraries (10–40%), with most of the imports instead being production-orientated. In contrast, newly added dependencies skew heavily toward testing and tooling, suggesting that when agents modify dependency manifests, they are mostly helping with project scaffolding rather than expanding runtime dependency graphs—an encouraging pattern from a security and maintenance perspective.
Second, language-specific choices follow ecosystem expectations. TypeScript imports focus on the front-end stack (react, next, lucide-react). Python shows a strong AI/data-science presence (crewai, openai, numpy, pandas) alongside ubiquitous HTTP clients (requests, httpx, aiohttp). Go leans toward infrastructure and cloud-native tooling (kubernetes packages, logging, CLI frameworks) alongside parser/protocol tools (tree-sitter, TypeScript compiler ports). C# imports include desktop/UI frameworks (Avalonia, Dock) and other Microsoft-ecosystem libraries, while new dependencies disproportionately configure test infrastructure (xUnit, NUnit, Moq). Overall, these trends show that agents generally choose sensible libraries for the given ecosystem.
RQ3 Summary
Across languages, agents import a broad set of production libraries but mostly add only test and developer tooling dependencies, indicating a conservative and tooling-oriented approach to dependency growth.
@@ 5. Threats to Validity
There are a few important limitations to consider when interpreting our findings. First, our results reflect behaviour in only a subset of AIDev, covering five deployed AI coding agents, four programming languages, and only high-quality repositories (100+ stars); other agents, languages, or repositories may exhibit different patterns. For internal validity, our extraction relies on regex parsing of file diffs, which can miss edge cases due to incomplete file context or complex syntax. Full abstract syntax tree parsing is not feasible because the diffs provide only partial files, so regex is the most practical option. We mitigated this threat by manually validating extraction from 100 diffs per language, and implementing comprehensive unit tests to ensure that expected syntax is correctly extracted.
@@ 6. Conclusion
Looking ahead, the next step is to probe not just which libraries agents use, but how well they use them. For example, do they reproduce known LLM failure modes, such as hallucinating library names (twistLibraryHallucinationsLLMs2025) or relying on deprecated features (wangLLMsMeetLibrary2025)? A deeper analysis of the code they generate would also help establish whether agents sensibly reuse existing libraries or quietly reimplement functionality. As agentic workflows continue to mature, understanding these behaviours will be key to integrating them safely and effectively into real software development practices.

### v2

@@ 4.2. Findings for RQ2 New Dependencies
For this RQ, we examine dependency manifests to investigate how often agents add new dependencies to projects. Results in Table 2.
Across all languages, dependency addition is extremely rare: only 1.3% of PRs introduce even a single new library. C# and Python are the only languages where dependency addition is non-negligible (6.1% and 2.8% of PRs). This conservative behaviour contrasts with typical human workflows in large ecosystems—such as TypeScript and Python—where adding a dependency to reuse existing functionality is common (Abdalkareem et al., 2020). Agents often struggle with installing packages and resolving environment conflicts (Arora et al., 2025), suggesting that their reluctance to add dependencies may reflect practical limitations rather than deliberate conservatism. Additionally, the presence of new dependencies is associated with a substantially lower merge rate only for Go PRs, where the merge rate drops by over 50%, compared to reductions of at most 10% in other languages.
When agents do introduce new dependencies, they generally do so with good versioning practices. Across all languages, 75.0% of new dependencies include an explicit version constraint, with TypeScript and Go specifying versions in every instance. This contrasts with previous, non-agentic, LLM research that observes version constraints mentioned in only $\approx$9% of library-related discussions (Raj and Costa, 2024). Whilst not directly comparable, this does indicate that agentic setups may support more robust dependency management over direct LLM usage; likely because agents benefit from full project context and direct interaction with dependency manifests.
Finally, we observe substantial diversity in the dependencies added: 1,119 unique libraries across 2,128 total additions. This high ratio of unique packages suggests that agents do not default to a small, preferred set of libraries when adding new dependencies. Much of this diversity comes from a small number of diverse PRs that create entirely new projects and therefore introduce many dependencies at once; a pattern consistent with known strengths of modern LLMs in scaffolding initial project structures (Rasnayaka et al., 2024).
RQ2 Summary. Agents rarely add new dependencies (1.3% of PRs), most commonly for C# (6.1%) and Python (2.8%). When they do add dependencies, they have strong version hygiene, specifying a constraint in 75.0% of cases.
@@ 4.3. Findings for RQ3 Choosing Libraries
For this RQ, we analyse trends in the libraries that agents most commonly use. Table 3 lists the most common libraries.
Across all languages, test frameworks and development-tooling libraries are widely used, but they make up only a minority of the distinct most-imported libraries (10–40%), with most imports instead being production-orientated. In contrast, newly added dependencies skew heavily toward testing and tooling, indicating that when agents modify dependency manifests, they prioritise scaffolding and validation over expanding runtime dependency graphs–a conservative and security-conscious pattern.
Library choices also align closely with language ecosystem norms. TypeScript imports focus on the front-end stack (react, next, lucide-react). Python shows a strong AI/data-science presence (crewai, openai, numpy, pandas) alongside various HTTP clients. Go leans toward infrastructure and cloud-native tooling (kubernetes packages, logging, CLI frameworks) alongside parser/protocol tools. C# imports include desktop/UI frameworks (Avalonia, Dock) other Microsoft-ecosystem libraries, while new dependencies disproportionately add test infrastructure (xUnit, NUnit, Moq). These results suggest that agentic workflows do not select libraries arbitrarily, but instead exhibit ecosystem-aware and risk-conscious behaviour when choosing both production libraries and new dependencies. RQ3 Summary. Across languages, agents import a broad set of production libraries but predominantly add testing and developer tooling dependencies, reflecting a conservative, ecosystem-aware approach to dependency growth.
@@ 5. Threats to Validity
First, our results reflect behaviour in only a subset of agent-authored PRs, covering five deployed AI coding agents, four programming languages, and repositories with 100+ stars; other agents, languages, or repositories may exhibit different patterns. For internal validity, our extraction relies on regex parsing of file diffs. This is the most practical choice because abstract syntax tree parsing is not feasible on partial diffs, but may miss edge cases due to incomplete context or complex syntax. We mitigate this threat through manual validation of 100 diffs per language and comprehensive unit testing.
@@ 6. Conclusion
Looking ahead, the next step is to probe not just which libraries agents use, but how well they use them. For example, do they reproduce known LLM failure modes, such as hallucinating library names (Twist et al., 2025b) or relying on deprecated features (Wang et al., 2025a)? A deeper analysis of the code they generate would also help establish whether agents sensibly reuse existing libraries or quietly reimplement functionality. As agentic workflows continue to mature, understanding these behaviours will be key to integrating them safely and effectively into real software development practices.
Acknowledgements.
This work was supported by the UKRI Centre for Doctoral Training in Safe and Trusted AI (EPSRC iCASE Award ref EP/Y528572/1).